1. You purchased 1000 shares of stock for $25 a share. After holding the stock for 6 years and not receiving any dividends, you sell the stock for $42 per share. What are the holding period and annual return on this investment?
a) 9.2% ; 1.63%
b) 18% ; 2.87%
c) 68% ; 9.03%
d) 62% ; 7.66%
Answer #1

c) 68% ; 9.03%

Holding period return = (P1-P0)/P0 Where,
= (42-25)/25 P0 = $          25
= 68% P1 = $          42
Holding period return = ((1+i)^n)-1 Where,
          0.68 = ((1+i)^6)-1 i = Annual return = ?
          1.68 = (1+i)^6 n = time = 6
          1.68 ^(1/6) = 1+i
1.090314 = 1+i
0.090314 = i
So,
Annual return is 9.03%